OpenStudy (anonymous):

WHAT IS THE CIRCUMFERENCE OF A CIRCLE IF THE RADIUS IS 21IN AND WHATS THE AREA OF THE CIRCLE IF THE RADIUS IS 21 IN?

OpenStudy (anonymous):

WHAT IS THE CIRCUMFERENCE OF A CIRCLE IF THE RADIUS IS 21IN AND WHATS THE AREA OF THE CIRCLE IF THE RADIUS IS 21

OpenStudy (anonymous):



OpenStudy (anonymous):

C=3.14x42...21+21=42....Circumference=131.88

OpenStudy (anonymous):

Circumference = 2 pi r = pi d. (because d=2r, you can always switch them) = 2 pi 21 = 138.88 in. Area = pi r^2 = pi 21^2 = 1384.74 in^2.

OpenStudy (anonymous):

THANKS!